OpenStudy (impracticaljoker):

9th grade math (01.07) Solve 3c + 4 = −20. (1 point) −5.3 6 8 −8

OpenStudy (impracticaljoker):

Isn't the answer 8?

OpenStudy (quickstudent):

This is actually really simple. You have to isolate c. Start like this: 3c + 4 = −20 3c + 4 - 4 = -20 - 4 ?

OpenStudy (quickstudent):

OK, let me check: 3c + 4 = -20 3c + 4 - 4 = -20 - 4 3c = -24 3c/3 = -24/3 c = -8

OpenStudy (impracticaljoker):

Oh, Forgot to keep it negative. Can you help me with another?

OpenStudy (quickstudent):

Sure! just open it in another question, okay?

OpenStudy (impracticaljoker):

k